OpenStudy (hba):

The area of the circle x^2+y^2-8x-24=0

hartnn (hartnn):

pi*r^2

OpenStudy (hba):

Thanks :)

OpenStudy (anonymous):

find r from the equation and apply \[\Pi r ^{2}\]

OpenStudy (anonymous):

What is the value of r here??

OpenStudy (hba):

\[r=2\sqrt{10}\]

OpenStudy (anonymous):

How ??

OpenStudy (hba):

\[40\pi \]

OpenStudy (anonymous):

the general equation of a cirle is \[(x-h)^{2} + (y-k)^{2} = r ^{2}\] compare this to your equation and find "r"

OpenStudy (hba):

\[r=\sqrt{g^2+f^2-c}\] \[r=\sqrt{16-(-24)}=\sqrt{16+24}=\sqrt{40}\]

OpenStudy (anonymous):

I know that.. I am asking find here proper value for r..

OpenStudy (hba):

r is r come on ?

OpenStudy (anonymous):

Are you not getting 8 ??

OpenStudy (anonymous):

Ok, I am wrong..

OpenStudy (anonymous):

Yes @hba you are right..

OpenStudy (hba):

@waterineyes Why do you keep confusing me lol haha

OpenStudy (hba):

So 40pi is final :D

OpenStudy (anonymous):

Ha ha ha ha..

hartnn (hartnn):

lol :P

hartnn (hartnn):

good going @waterineyes

OpenStudy (anonymous):

in this case compare your equation to \[x ^{2} + y ^{2} + 2gx + 2fy + c = 0\] where \[r = \sqrt{g ^{2} + f ^{2} - c}\]

OpenStudy (anonymous):

Thanks @hartnn

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

@tomiko you are late dear but it is okay.. ha ha ha ha..

OpenStudy (anonymous):

Anyways, thanks for reminding me the equation of circle..

OpenStudy (hba):

@tomiko No i do not get it bcx i have already got it :D

OpenStudy (hba):

@waterineyes Don't worry :)

OpenStudy (anonymous):

Notification came first and then your comment.. Great..

OpenStudy (hba):

You should buy a usb for your laptop

OpenStudy (anonymous):

For what??

OpenStudy (anonymous):

And I am having a pc not laptop..

OpenStudy (hba):

Like wifi

OpenStudy (anonymous):

haha no problem. you guys are funny! =))

OpenStudy (hba):

Still counts...

OpenStudy (hba):

No we do do not work in the circus,we are just havin fun ;D @tomiko

OpenStudy (anonymous):

I am comfortable with this slow running and at the same time ruining tortoise Internet Stick..

OpenStudy (hba):

The center of the circle x^2+y^2-6x+8y-24=0 I got (3,-4) right or wrong ? @waterineyes

OpenStudy (hba):

@waterineyes How can you live in this world with such an awesome internet connection :)

OpenStudy (anonymous):

It is awesome that is why I am living and at the same time leaving on it..

OpenStudy (hba):

@waterineyes What about my ques ?

OpenStudy (anonymous):

(3,-4) is right..

OpenStudy (hba):

So none of the above :D

OpenStudy (anonymous):

No..

OpenStudy (hba):

Shuuu

OpenStudy (hba):

It is not in options lol

OpenStudy (anonymous):

Data is inadequate..

OpenStudy (hba):

For you ^

OpenStudy (anonymous):

Yep.. Going to sleep.. 12:51 am here.. Good Night. @hba

OpenStudy (hba):

Please see my last quessss :(

OpenStudy (hba):

@waterineyes ??